OpenStudy (anonymous):

Write the equation of the line with a slope of -5 and a y-intercept of (0,3). Write the equation of the line with a slope of -1/3 and passing through the point (6, -4). Write the equation of the line passing through the points (0, -4) and (-2, 2). Write the equation of the line passing through the points (-6,1) and (-4,2). Write the equation of the line with an undefined slope, passing through the point (2, 5).

OpenStudy (austinl):

Whoa, that's a lot of questions.... I suppose lets get crackin'.

OpenStudy (anonymous):

I think I got the first one.

OpenStudy (austinl):

Lets start with the first one. \(Write ~the ~equation ~of ~the ~line ~with ~a ~slope ~of -5~ and ~a ~y-intercept ~of ~(0,3).\) The equation of a line in slope intercept form is, \(y=mx+b\) \(m=slope\) \(b=y-intercept\) What do you get as an answer for the first question?

OpenStudy (anonymous):

y=3-5x

OpenStudy (austinl):

Yep, that is correct! Now onto the second question.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Is it y=-x/3-2 ?

OpenStudy (austinl):

It is indeed!

OpenStudy (anonymous):

Yay! I think I actually do get it now.

OpenStudy (austinl):

Do you know how to do the next one?

OpenStudy (anonymous):

Yeah I think so, I'll try it and give you my answer.

OpenStudy (anonymous):

y=-3x-4 ?

OpenStudy (austinl):

Correctomundo.

OpenStudy (anonymous):

Thank you. I get it now. c:

OpenStudy (austinl):

No problem!